OpenStudy (anonymous):

Evaluate \[\sum_{n=3}^{7}~{n\over4}\]

OpenStudy (anonymous):

i got \(0\)...?

OpenStudy (harsimran_hs4):

no zero is perhaps not correct answer... \[\frac{ 1 }{ 4 } \sum_{3}^{7} n \] can you do it now?

OpenStudy (anonymous):

\[\sum_{n=3}^7\frac{n}{4}=\frac{3}{4}+\frac{4}{4}+\frac{5}{4}+\frac{6}{4}+\frac{7}{4}\]

OpenStudy (anonymous):

oh i posted the wrong one...that one is \(6{1\over4}\) i meant to post this:\[\large\sum_{n=2}^{8}-(n-5)\]

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

plug and chug

OpenStudy (anonymous):

yea i got \(0\), is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[[-(2-5)]+[-(3-5)]+[-(4-5)]+[-(5-5)]+[-(6-5)]+[-(7-5)]+[-(8-5)]\]\[(3)+(2)+(1)+(0)+(-1)+(-2)+(-3)=0\]

OpenStudy (anonymous):

thanks :)